OpenStudy (anonymous):

If F(x) = x - 1, which of the following is the inverse of F(x)? A. F-1(x) = 1 - x B. F-1(x) = x - 1 C. F-1(x) = x + 1 D. F-1(x) = x

OpenStudy (anonymous):

\[f(x)=x-1\]\[Let~y=x-1\]\[y+1=x\]\[f^{-1}(x)=y-1\]\[Therefore,f^{-1}(x)=x-1\]

OpenStudy (anonymous):

which option is correct? @ShyannaD

OpenStudy (anonymous):

i am not for sure

OpenStudy (anonymous):

Answer is B